The Bonavista Patient Transfer Company is considering the purchase of a new ambulance. The decision will rest partly on the anticipated distance to be driven next year. The kilometers driven during the past five years are as follows: 

Year 

1 3000 

2 4000 

3 3400 

4 3800 

5 3700 

.1. Forecast the number of kilometers for next year using a two-year moving average. 

2. Find the MAD based on the two-year moving average forecast in part (a). (Hint: You will have only three years of matched data.) 

3. Use a weighted two-year moving average with weights of 0.4 and 0.6 to forecast next year's mileage. (The weight of 0.6 is for the most recent year.) 

4. What MAD results from using this approach to forecasting? (Hint: You will have only three years of matched data.)
Answer #1

ANSWER::

1).

Forecast = Sigma(Demand in previous N periods) / N

Where N = 2

FORECAST 3 = (3000 + 4000) / 2 = 3500

FORECAST 4 = (4000 + 3400) / 2 = 3700

FORECAST 5 = (3400 + 3800) / 2 = 3600

FORECAST 6 = (3800 + 3700) / 2 = 3750

2).

MAD

PERIOD

ACTUAL

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

3000

2

4000

3

3400

3500

-100

100

4

3800

3700

100

100

5

3700

3600

100

100

SIGMA

100

300

MAD = SIGMA(ABS DEV) / N

MAD = 300 / 3 = 100

3) and 4)

. Forecast = Sigma(weight for period N * Demand for period N) / Sigma(weights)

Where the highest weights are multiplied with the most recent demand value.

FORECAST 3 = (0.4 * 3000) + 0.6 * 4000) / (0.4 + 0.6) = 3600

FORECAST 4 = (0.4 * 4000) + 0.6 * 3400) / (0.4 + 0.6) = 3640

FORECAST 5 = (0.4 * 3400) + 0.6 * 3800) / (0.4 + 0.6) = 3640

FORECAST 6 = (0.4 * 3800) + 0.6 * 3700) / (0.4 + 0.6) = 3740

MAD

PERIOD

ACTUAL

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

3000

2

4000

3

3400

3600

-200

200

4

3800

3640

160

160

5

3700

3640

60

60

SIGMA

20

420

MAD = SIGMA(ABS DEV) / N

MAD = 420 / 3

MAD= 140